Initially the ball is dropped from height h above the ground and the coefficient of Restitution is 0.9 count the no of bounces before the ball stops bouncing

Grade:12

1 Answers

Balaji
6 years ago
e=√(hn/h)Given e =0.9(0.9)^2=hn/hi.e hn=(0.9)^2*hFor renouncehn=(e)^2n*h(0.9)^2*h=(e)^2n*hWhen bases are same power should be same i.e 2n=2Then n=1